OpenStudy (anonymous):

Find the geometric mean between 25 and 144

OpenStudy (anonymous):

60

OpenStudy (anonymous):

its basically \[\sqrt{25*144}\Rightarrow\sqrt{25}*\sqrt{144}\Rightarrow5*12=60\]

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

yes good good